# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ 5.1.3 Challenges for Non-differentiable Neural Networks
Recall that Theorem 5.1 holds for neural networks with locally Lipschitz gradient, and thus excludes ReLU neural networks. We first provide an empirical evidence that early directional convergence also occurs in ReLU networks, and then briefly describe the challenges in extending Theorem 5.1 to ReLU networks. Let $\sigma(x)=\max(x,0)$. We train the following $3-$homogeneous ReLU network $\mathcal{H}(\mathbf{x};\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{W}_{3})=\mathbf{1}^{\top}\sigma(\mathbf{W}_{3}\sigma(\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x})))$, where $\mathbf{W}_{1}\in\mathbb{R}^{20\times 10}$, $\mathbf{W}_{2},\mathbf{W}_{3}\in\mathbb{R}^{20\times 20}$ and $\mathbf{1}\in\mathbb{R}^{20}$ is the all-one vector. For training, we minimize the square loss with respect to the output of a smaller neural network *$\mathcal{H}^{*}(\mathbf{x};\mathbf{W}_{1}^{*},\mathbf{W}_{2}^{*},\mathbf{W}_{3}^{*})=\mathbf{1}^{\top}\sigma(\mathbf{W}_{3}^{*}\sigma(\mathbf{W}_{2}^{*}\sigma(\mathbf{W}_{1}^{*}\mathbf{x})))$, where the entries of of $\mathbf{W}_{1}^{*}\in\mathbb{R}^{2\times 10}$, $\mathbf{W}_{2}^{*}\in\mathbb{R}^{3\times 2},\mathbf{W}_{3}^{*}\in\mathbb{R}^{3\times 3}$ are drawn from standard normal distribution and $\mathbf{1}\in\mathbb{R}^{3}$ is the all-one vector. The training data has 100 points sampled uniformly from unit sphere in $\mathbb{R}^{10}$. We define $\mathbf{w}$ and $\tilde{\mathbf{w}}$ in a similar way as in the previous experiment, and $\mathcal{N}(\mathbf{w})$ denotes the NCF for this case. The network is trained for 80000 iterations using gradient descent with step-size $10^{-3}$, and the initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, where $\delta=0.01$ and $\mathbf{w}_{0}$ is a random unit norm vector. From Figure 3(a), we observe that the training loss does not change much, and the norm of the weights increases by a multiplicative factor but remains small. Also, from Figure 3(b), it is clear that the weights approximately converge in direction to a KKT point of the constrained NCF. Furthermore, similar to the previous experiment, low-rank structure emerges among the weights of neural networks; see Figure 5.
Figure 4: Early training dynamics of a ReLU neural network. Panel $(a)$: the evolution of training loss (normalized with respect to loss at initialization) and the $\ell_{2}$-norm of all the weights with iterations. Panel $(b)$: the evolution of $\partial\mathcal{N}(\tilde{\mathbf{w}}(t))^{\top}\tilde{\mathbf{w}}(t)/\|\partial\mathcal{N}(\tilde{\mathbf{w}}(t))\|_{2}$ (a measure of directional convergence of $\mathbf{w}(t)$). The loss does not change much, and the norm of weights has increased by a constant factor but remains small. Also, the weights approximately converge in direction to a KKT point of the constrained NCF.
We next describe the difficulty in extending Theorem 5.1 to ReLU networks. Consider a neural network that satisfies 1, except that the gradient is not assumed to be locally Lipschitz (this will include ReLU networks). Then, assuming square loss, we can define the gradient flow dynamics using the Clarke subdifferential as follows
It is also well-known that the Clarke differential of an $L$-homogeneous function is $(L-1)$-homogeneous [32, Theorem B.2]. Thus, if we define $\mathbf{s}(t)=\frac{1}{\delta}\mathbf{w}\left(\frac{t}{\delta^{L-2}}\right)$, then similar to (8), we can show that $\mathbf{s}(t)$ will evolve according to
Following the proof sketch of Theorem 5.1 described in Section 5.1.1, to show approximate directional converegence of $\mathbf{w}(t)$, we would need to first show that for small enough $\delta$, the solutions of (12) and (11) will be sufficiently small for sufficiently long time. From comparing (12) and (11), it is clear that for small $\delta$, the two solutions will be close in the beginning. However, we can not claim that the solutions will remain close for sufficiently long time by choosing small enough $\delta$. In fact, the proof of the analogous result in the context of Theorem 5.1 crucially relied on gradients being locally Lipschitz, which is not the case here. We believe that proving this would be a crucial step towards establishing directional convergence for ReLU networks, and defer it for future investigations.
Figure 5: Panel $(a)$: $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at iteration 80000. Panel $(b)$: $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at iteration 80000. Panel $(c)$: $\mathbf{W}_{3}/\|\mathbf{w}\|_{2}$ at iteration 80000. Each curve in the plots correspond to a row of the corresponding weights matrix.
5.2 Directional Convergence Near Saddle Points
For small initializations and square loss, several recent studies have also observed a so-called saddle-to-saddle dynamics during training of neural networks; see, e.g., [34, 35], where it is observed that during training the weights of neural networks passes through a sequence of saddle points. Moreover, recent theoretical results have further revealed that only a certain number of weights are non-zero at saddle point encountered during training [8, 35]. Motivated by these observations, in this section we analyze the gradient flow dynamics near certain saddle points, specifically those saddle points where only a certain number of weights are non-zero while the others are zero. It is worth noting that [1] also analyzed the gradient flow dynamics near similar types of saddle points for two-homogeneous neural networks. In this section, we assume that the neural network has a separable structure. Specifically, we assume that the weights of the neural network $\mathbf{w}$ can be partitioned into two sets, that is, $\mathbf{w}=\left[\mathbf{w}_{n},\mathbf{w}_{z}\right]$, such that the output of the neural network can be written as
$\mathcal{H}(\mathbf{x};\mathbf{w})=\mathcal{H}_{n}(\mathbf{x};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z}),$
where $\mathcal{H}_{n}(\mathbf{x};\mathbf{w}_{n})$ and $\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z})$ are assumed to satisfy 1. Now, let $\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})$ and $\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})$ be defined similarly to $\mathcal{H}(\mathbf{X};\mathbf{w})$, then for square loss, the training loss is

### v2

@@ 4.1.2 A Closer Look at the Weights
We train the following $3-$homogeneous neural network $\mathcal{H}(\mathbf{x};\mathbf{v},\mathbf{U})=\mathbf{v}^{\top}\max(\mathbf{U}\mathbf{x},\mathbf{0})^{2}$, where $\mathbf{v}\in\mathbb{R}^{20\times 1}$ and $\mathbf{U}\in\mathbb{R}^{20\times 10}$, by minimizing the square loss with respect to the output of a smaller neural network $\mathcal{H}^{*}(\mathbf{x};\mathbf{v}_{*},\mathbf{U}_{*})=\mathbf{v}_{*}^{\top}\max(\mathbf{U}_{*}\mathbf{x},\mathbf{0})^{2}$, where the entries of $\mathbf{v}_{*}\in\mathbb{R}^{2\times 1}$, $\mathbf{U}_{*}\in\mathbb{R}^{2\times 10}$ are drawn from the standard normal distribution. The training data has 100 points sampled uniformly from unit sphere in $\mathbb{R}^{10}$. Let $\mathbf{w}$ be a vector containing all the entries of $\mathbf{v}$ and $\mathbf{U}$, and $\widetilde{\mathbf{w}}=\mathbf{w}/\|\mathbf{w}\|_{2}.$ We train for 50360 iterations using gradient descent with step-size $2\cdot 10^{-2}$, and the initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, where $\delta=0.05$ and $\mathbf{w}_{0}$ is a random unit norm vector. Figure 1 depicts the evolution of training loss and $\ell_{2}-$norm of the weights. Note that the loss almost remains the same, and the norm of the weights has increased but only by a constant factor and is still small. A unit norm vector is a KKT point of the constrained NCF if the vector and the gradient of the NCF at that vector are parallel (see Lemma 11). Hence, to measure the directional convergence of $\mathbf{w}$ to the KKT point of the constrained NCF, in Figure 1 we plot the evolution of the inner product between $\widetilde{\mathbf{w}}(t)$ and $\nabla\mathcal{N}(\widetilde{\mathbf{w}}(t))/\|\nabla\mathcal{N}(\widetilde{\mathbf{w}}(t))\|_{2}$, where $\mathcal{N}(\cdot)$ denotes the NCF for this example. Clearly, in accordance with Theorem 1, the weights approximately converge in direction to a KKT point of the constrained NCF.
We next take a closer look at the weights of the neural networks at iteration 50360, i.e., when the weights have approximately converged in direction to a KKT point of the constrained NCF. In Figure 2, we plot the normalized absolute value of the weights at initialization and at iteration 50360. At initialization the weights appear to be random, as expected, but at iteration 50360, they are more structured. The hidden weights (${\mathbf{U}}$) appear to be of rank-one as only one row is non-zero, and the outer weight (${\mathbf{v}}$) has a single non-zero entry. This suggests that weights converged towards a KKT point of the constrained NCF that have low-rank hidden weights. Similar behavior is observed with deeper networks and other activation functions as well (see Section 5.3), and has also been observed in some previous works (Zhou et al., 2022; Atanasov et al., 2022).
Since the constrained NCF is a non-convex problem with potentially multiple KKT points, an important question arising from the above observation is why gradient descent converges towards low-rank KKT points. A more basic question, addressed in the next section, is whether low-rank KKT points always exist for the constrained NCF, and can they be characterized? Given the complex nature of the problem, their existence is not immediately clear, especially for deep neural networks. We next study the KKT points of the constrained NCF for feed-forward neural networks with Leaky ReLU and polynomial Leaky ReLU activation functions. For arbitrary training data, necessary and sufficient conditions are derived for a set of weights to be a KKT point of the NCF that have rank-one hidden weights.
(a) At initialization
(b) At iteration 50360
Figure 2: Normalized absolute value of the weights at different stages of training.
5 KKT Points of Constrained NCF
This section presents necessary and sufficient conditions for a set of weights to be a KKT point of the constrained NCF of feed-forward neural network that have rank-one hidden weights. For a given input ${\mathbf{x}}$, the output of an $L$-layer feed-forward neural network $\mathcal{H}$ is defined as
$\mathcal{H}({\mathbf{x}};{\mathbf{W}}_{1},\cdots,{\mathbf{W}}_{L})={\mathbf{W}}_{L}\sigma\left({\mathbf{W}}_{L-1}\cdots\sigma\left({\mathbf{W}}_{1}{\mathbf{x}}\right)\cdots\right)\in{\mathbb{R}},$ (9)
where ${\mathbf{W}}_{l}\in{\mathbb{R}}^{k_{l}\times{k_{l-1}}}$ are the trainable weights, and $k_{0}=d$ and $k_{L}=1$. The activation function $\sigma(\cdot)$ is applied coordinate-wise to vectors, and in this section is assumed to be of the form $\sigma(x)=\max(\alpha x,x)^{p}$, where $p\in{\mathbb{N}}$ and $\alpha\in{\mathbb{R}}$, which includes deep ReLU and polynomial ReLU neural networks. The corresponding constrained NCF is the following optimization problem:
$\max_{{\mathbf{W}}_{1},\cdots,{\mathbf{W}}_{L}}\mathcal{N}({\mathbf{W}}_{1},\cdots,{\mathbf{W}}_{L}):={\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};{\mathbf{W}}_{1},\cdots,{\mathbf{W}}_{L}),\text{ s.t. }\|{\mathbf{W}}_{1}\|_{F}^{2}+\cdots+\|{\mathbf{W}}_{L}\|_{F}^{2}=1.$ (10)
Note that, we have considered the NCF in the above equation with respect to ${\mathbf{y}}$ for convenience, it may be replaced with any other vector.
5.1 Leaky ReLU
We first consider the activation function of the form $\sigma(x)=\max(x,\alpha x)$, for some $\alpha\in{\mathbb{R}}$.
Theorem 3.
Let $\mathcal{H}$ be an $L-$layer feed-forward neural network as defined in eq. 9, where $L\geq 2$ and $\sigma(x)=\max(x,\alpha x)$ for some $\alpha\in{\mathbb{R}}$. Let $\overline{{\mathbf{W}}}_{1:L}$ be defined as
$\overline{{\mathbf{W}}}_{1:L}:=(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L-1},\overline{{\mathbf{W}}}_{L})=({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top},\cdots,{\mathbf{a}}_{L-1}{\mathbf{b}}_{L-1}^{\top},\overline{{\mathbf{w}}}^{\top}),\mbox{ where $\|{\mathbf{a}}_{l}\|_{2}=\|{\mathbf{b}}_{l}\|_{2},$ for all $l\in[L-1].$}$
If $\overline{{\mathbf{W}}}_{1:L}$ is a KKT point of eq. 10, then $\|{\mathbf{a}}_{l}\|_{2}^{2}=1/\sqrt{L}$, for all $l\in[L-1],$ and $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/L$.
  • Necessary and sufficient condition. Suppose $\alpha\neq 1$ and the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: ${\mathbf{a}}_{l}={\mathbf{b}}_{l+1}$, for all $l\in[L-2]$, and $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, for some $q\in\{-1,1\}$. Further, ${\mathbf{b}}_{1}$ is a non-zero KKT point of
    $\max_{{\mathbf{u}}}\sum_{i=1}^{n}y_{i}\sigma^{L-1}({\mathbf{x}}_{i}^{\top}{\mathbf{u}}),\text{ such that }\|{\mathbf{u}}\|_{2}^{2}=1/\sqrt{L}.$ (11)
  • Necessary and sufficient condition. Suppose $\alpha=1$, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: $q_{l}{\mathbf{a}}_{l}={\mathbf{b}}_{l+1},$ for all $l\in[L-2]$, and $q_{L-1}{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, where $q_{l}\in\{-1,1\}$. Further, ${\mathbf{b}}_{1}$ is a non-zero KKT point of
    $\max_{{\mathbf{u}}}\sum_{i=1}^{n}y_{i}\sigma^{L-1}({\mathbf{x}}_{i}^{\top}{\mathbf{u}}),\text{ such that }\|{\mathbf{u}}\|_{2}^{2}=1/\sqrt{L}.$ (12)
In the above theorem, $\{{\mathbf{a}}_{i},{\mathbf{b}}_{i}\}_{i=1}^{L-1}$ are vectors used to build the rank-one hidden weights. Our first result provides the norm of these vectors. Next, for $\alpha\neq 1$ and assuming the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, we present necessary and sufficient conditions, which can be divided in two parts. First is the alignment: ${\mathbf{a}}_{l}={\mathbf{b}}_{l+1}$, for all $l\in[L-2]$, and $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, for some $q\in\{-1,1\}$, which implies that the column-space of $\overline{{\mathbf{W}}}_{l}$ and row-space of $\overline{{\mathbf{W}}}_{l+1}$ must be aligned. The second part requires ${\mathbf{b}}_{1}$ to be a KKT point of eq. 11, a smaller optimization problem, where note that $\sigma^{L-1}(\cdot)$ denotes the function resulting from composing $\sigma(\cdot)$ with itself $(L-1)$ times. The results for $\alpha=1$ are mostly similar, but do not require any non-negativity assumption.
5.2 Polynomial Leaky ReLU
We next provide necessary and sufficient conditions for activation function of the form $\sigma(x)=\max(x,\alpha x)^{p}$, for some $\alpha\in{\mathbb{R}}$, $p\geq 2$ and $p\in{\mathbb{N}}$. In the following theorem, for a given $p$, we define $\hat{p}\coloneqq p^{L-1}+p^{L-2}+\cdots+1.$
Theorem 4.
Let $\mathcal{H}$ be an $L-$layer feed-forward neural network as defined in eq. 9, where $L\geq 2$ and $\sigma(x)=\max(x,\alpha x)^{p}$ for some $\alpha\in{\mathbb{R}}$, $p\geq 2$ and $p\in{\mathbb{N}}$. Let $\overline{{\mathbf{W}}}_{1:L}$ be defined as
$\overline{{\mathbf{W}}}_{1:L}:=(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L-1},\overline{{\mathbf{W}}}_{L})=({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top},\cdots,{\mathbf{a}}_{L-1}{\mathbf{b}}_{L-1}^{\top},\overline{{\mathbf{w}}}^{\top}),\mbox{ where $\|{\mathbf{a}}_{l}\|_{2}=\|{\mathbf{b}}_{l}\|_{2},$ for all $l\in[L-1].$}$
If $\overline{{\mathbf{W}}}_{1:L}$ is a KKT point of eq. 10, then $\|{\mathbf{a}}_{l}\|_{2}^{4}=p^{L-l}/\hat{p}$, for all $l\in[L-1],$ and $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/\hat{p}$.
  • Necessary and sufficient condition. Suppose $\alpha\neq 1$ and the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: ${\mathbf{b}}_{l}={\mathbf{a}}_{l-1}/p^{1/4}$, for $12\leq l\leq L-$ 1, and $\overline{{\mathbf{w}}}=q{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, for some $q\in\{-1,1\}$. Further, for $11\leq l\leq L-$ 1, each non-zero entry of ${\mathbf{a}}_{l}$ has identical values, and ${\mathbf{b}}_{1}$ is a non-zero KKT point of
    $\max_{{\mathbf{u}}}\sum_{i=1}^{n}y_{i}\sigma^{L-1}({\mathbf{x}}_{i}^{\top}{\mathbf{u}}),\text{ such that }\|{\mathbf{u}}\|_{2}^{2}=\sqrt{p^{L-1}/\hat{p}}.$ (13)
  • Necessary and sufficient condition. Suppose $\alpha=1$, then $\overline{{\mathbf{W}}}_{1:L}$ is a non-zero KKT point of eq. 10 if and only if the following holds: ${\mathbf{b}}_{l}=q_{l}{\mathbf{a}}_{l-1}/p^{1/4},\overline{{\mathbf{w}}}=q_{L}{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, if $p$ is odd, and ${\mathbf{b}}_{l}=q_{l}|{\mathbf{a}}_{l-1}|/p^{1/4},\overline{{\mathbf{w}}}=q_{L}|{\mathbf{a}}_{L-1}|/(p\hat{p})^{1/4}$, if $p$ is even, where $q_{l}\in\{-1,1\}$, for all $2\leq l\leq L$. Further, for $11\leq l\leq L-$ 1, each non-zero entry of ${\mathbf{a}}_{l}$ has identical absolute values, and ${\mathbf{b}}_{1}$ is a non-zero KKT point of
    $\max_{{\mathbf{u}}}\sum_{i=1}^{n}y_{i}\sigma^{L-1}({\mathbf{x}}_{i}^{\top}{\mathbf{u}}),\text{ such that }\|{\mathbf{u}}\|_{2}^{2}=\sqrt{p^{L-1}/\hat{p}}.$ (14)
Overall, the results are mostly analogous to Theorem 3, with one key difference: the non-zero entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1}$ must have identical (absolute) values. This difference is also evident in Figure 3, which depicts KKT points of 3-layer neural network with ReLU and squared-ReLU activation. For ReLU, in contrast with squared-ReLU, multiple rows and columns of the weights are non-zero. Additionally, the KKT point exhibits sparsity, with ReLU activation having multiple rows and columns with zero norm, while squared-ReLU activation has only one non-zero row and column. We observe this difference consistently in our experiments.
(a) ReLU activation
(b) Squared-ReLU activation
Figure 3: Absolute value of a KKT point $\left(\overline{{\mathbf{W}}}_{1},\overline{{\mathbf{W}}}_{2},\overline{{\mathbf{W}}}_{3}\right)$ of eq. 10 for ReLU activation $(\max(x,0))$ and squared-ReLU activation $(\max(x,0)^{2})$, where $L=3$, $n=100$ and ${\mathbf{x}}_{i}\sim_{\text{i.i.d}}\mathcal{N}(0,{\mathbf{I}}_{10}),y_{i}\sim_{\text{i.i.d}}\mathcal{N}(0,1)$. For ReLU, the non-zero rows/columns are scalar multiple of each other, implying the hidden weights have rank one. For squared-ReLU, it is clear that the hidden weights have rank one.
The results of Theorem 3 and 4 contain sufficient conditions which imply that KKT points of the constrained NCF with rank-one hidden weights always exists for arbitrary training data or depth of the neural network. The sufficient conditions also provide an approach to obtain a KKT point of the constrained NCF using KKT points of a smaller optimization problem. Additionally, they also contain matching necessary conditions, where, for $\alpha\neq 1$, we had to further assume that $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ have non-negative entries. This implies that if, for $\alpha\neq 1$, there is a KKT point with rank-one hidden weights such that $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ have non-negative entries, then it must be of the form described in the above theorems. Interestingly, in our experiments with random training data (Section 5.3), we found that the KKT points found via gradient-based method generally satisfy this non-negativity assumption, along with having rank-one hidden weights.
Remark 1.
In the above theorems, we focus on non-zero KKT points, because zero KKT points may not have a rich structure. For instance, using Lemma 11, it can be shown that if the network’s output is always zero, the corresponding weights are zero KKT points. For ReLU activation, the output is zero if the input is non-positive. Thus, for an $L$-layer ReLU network, the output will be zero if any matrix in $({\mathbf{W}}_{2},\dots,{\mathbf{W}}_{L-1})$ has non-positive entries, making those weights zero KKT points.
5.3 Numerical Experiments
In the above theorems, we have made two main assumptions on the KKT points of the NCF $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})$:
  1. 1.
    Rank-one hidden weights. $\overline{{\mathbf{W}}}_{l}={\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top},$ where $\|{\mathbf{a}}_{l}\|_{2}=\|{\mathbf{b}}_{l}\|_{2},$ for all $l\in[L-1].$
  2. 2.
    Non-negative weights. If $\sigma(x)=\max(x,\alpha x)^{p}$ and $\alpha\neq 1$, then $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative.
To validate these assumptions, we conduct numerical experiments for two and three layer neural network with activation function $\sigma(x)=\max(\alpha x,x)^{p}$, where $p\in\{1,2\}$ and $\alpha\in\{0,0.1,1\}$. For our experiments, we generate a random set of training data, and then obtain a KKT point of the corresponding constrained NCF by maximizing the NCF via gradient ascent with random initialization; [^3] Since gradient ascent can become unbounded, we use projected gradient ascent, i.e., after every gradient ascent step we normalize the weights to have unit norm. However, it also turns out that for homogeneous objectives, projected gradient ascent with fixed step-size is equivalent to gradient ascent with adaptive step-size, up to a scaling factor (see Lemma 19). more details are provided in the Section C.4. The results presented below are for 30 such random instances. Also, as discussed earlier, the gradient flow of the training loss converges to the same KKT point as the gradient flow of the NCF. Thus, using gradient ascent to find the KKT point of the constrained NCF also provides insights into the KKT point towards which gradient flow of the training loss converges in the early stages of training.
Rank-one assumption: For any set of matrices $\{{\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p}\}$, we use
$\kappa({\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p})=:\max_{1\leq i\leq p}\left(1-\frac{\|{\mathbf{Z}}_{i}\|_{2}}{\|{\mathbf{Z}}_{i}\|_{F}}\right)$
to measure how close are each of $\{{\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p}\}$ to being rank-one. If $\kappa({\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p})=0$, then each of $\{{\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p}\}$ are rank one, and thus, for smaller $\kappa({\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p})$, they are approximately rank one. Also, $\kappa(\cdot)$ is related to the notion of numerical rank (Rudelson & Vershynin, 2007), a stable relaxation of the rank. Now, from the values in Table 1, we observe that for both two-layer and three-layer neural network, and with different activation functions, each of the hidden weights in the KKT point obtained via gradient ascent are approximately of rank one in all the instances considered (since we are reporting the maximum value of $\kappa(\cdot)$ across all the random instances.)
Non-negativity assumption: For any set $\{{\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p}\}$ containing matrices and/or vectors, we use
$\rho({\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p})=:\max_{1\leq i\leq p}\left(\frac{\|\max(0,-{\mathbf{Z}}_{i})\|_{F}}{\|{\mathbf{Z}}_{i}\|_{F}}\right)$
to measure how close are each of $\{{\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p}\}$ to being non-negative. Note that, if ${\mathbf{Z}}_{i}$ is a vector, then $\|{\mathbf{Z}}_{i}\|_{F}$ denotes its $\ell_{2}$-norm. If $\rho({\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p})=0$, then each of $\{{\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p}\}$ are non-negative, and thus, for smaller $\rho({\mathbf{Z}}_{1},\cdots,{\mathbf{Z}}_{p})$, they are approximately non-negative. From the values in Table 1, we observe that for both two-layer and three-layer neural network, and with different activation functions, the appropriate set of weights in the KKT point obtained via gradient ascent are approximately non-negative in all the instances considered. Note that, we did not consider linear or monomial activation $(\sigma(x)=x^{p})$ since in those cases we did not make any non-negativity assumption.
We conclude this section by again noting that under the rank-one and non-negativity assumption considered here, the sufficient conditions of Theorem 3 and 4 are also necessary conditions, and thus, the KKT points must be of the form described in those results. Since the rank-one and non-negativity assumption is satisfied for all the instances considered, it implies the KKT points are of form described in Theorem 3 and 4 for all the instances considered. Also, note that the constrained NCF is a non-convex problem and could have KKT points that don’t satisfy these assumptions, but gradient ascent seems to avoid them. Understanding this phenomenon is an interesting future direction, which would require analyzing the dynamics of gradient ascent.
$\alpha=0$ $\alpha=0.1$ $\alpha=1$
$p=1$ $5.93\cdot 10^{-9}$ $1.79\cdot 10^{-8}$ $2.07\cdot 10^{-8}$
$p=2$ $1.76\cdot 10^{-10}$ $1.67\cdot 10^{-10}$ $6.86\cdot 10^{-11}$
Three-layer
$\alpha=0$ $\alpha=0.1$ $\alpha=1$
$p=1$ $2.97\cdot 10^{-12}$ $1.22\cdot 10^{-11}$ $1.9\cdot 10^{-10}$
$p=2$ $1.48\cdot 10^{-11}$ $1.12\cdot 10^{-11}$ $7.93\cdot 10^{-11}$
Two-layer
Table 1: The left table contains the maximum value of $\kappa(\overline{{\mathbf{W}}}_{1},\overline{{\mathbf{W}}}_{2})$ across 30 different random instances for a three-layer neural network with activation function $\max(x,\alpha x)^{p}$, where $(\overline{{\mathbf{W}}}_{1},\overline{{\mathbf{W}}}_{2},\overline{{\mathbf{W}}}_{3})$ is the KKT point of the constrained NCF obtained via gradient ascent. The right table is similar but for two-layer neural network.
$\alpha=0$ $\alpha=0.1$
$p=1$ $4.44\cdot 10^{-16}$ $3.28\cdot 10^{-16}$
$p=2$ $5.16\cdot 10^{-6}$ $5\cdot 10^{-6}$
Three-layer
$\alpha=0$ $\alpha=0.1$
$p=1$ $1.82\cdot 10^{-7}$ $2.77\cdot 10^{-7}$
$p=2$ $3.95\cdot 10^{-7}$ $3.58\cdot 10^{-7}$
Two-layer
Table 2: Suppose $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})$ is the KKT point of the constrained NCF obtained via gradient ascent for an $L-$layer neural network, and let ${\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top}$ be the rank-one approximation of ${\mathbf{W}}_{l}$ such that $\|{\mathbf{a}}_{l}\|_{2}=\|{\mathbf{b}}_{l}\|_{2}$, for all $l\in[L-1]$. The left (right) table contains the maximum value of $\rho({\mathbf{a}}_{1},{\mathbf{a}}_{2}{\mathbf{b}}_{2}^{\top},\cdots{\mathbf{a}}_{L-1}{\mathbf{b}}_{L-1}^{\top})$ across 30 different random instances for a three-layer (two-layer) neural network with activation function $\max(x,\alpha x)^{p}$. Note that if ${\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top}$ is non-negative, then we can choose ${\mathbf{a}}_{l}$ and ${\mathbf{b}}_{l}$ such that they are also non-negative. Hence, showing ${\mathbf{a}}_{1},{\mathbf{a}}_{2}{\mathbf{b}}_{2}^{\top},\cdots{\mathbf{a}}_{L-1}{\mathbf{b}}_{L-1}^{\top}$ are non-negative is equivalent to showing $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative.
5.4 Proof Overview
We now provide a brief overview of the proof technique for the above theorems; the details can be found in Appendix C. We start with the following lemma, which is important for our proof.
Lemma 5.
Let ${\mathcal{H}}$ be an $L$-layer feed-forward neural network whose output is as defined in eq. 9, where $L\geq 2$. Suppose $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})$ is a non-zero KKT points of the constrained NCF in eq. 10. If $\sigma(x)=\max(\alpha x,x)^{p}$, for some $p\in{\mathbb{N}}$ and $\alpha\in{\mathbb{R}}$, then
$\text{diag}\left(\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}\right)=p\cdot\text{diag}\left(\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1}\right),\text{ for all }l\in[L-1],$ (15)
and, if $\sigma(x)=x$, then
$\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}=\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1},\text{ for all }l\in[L-1].$
Furthermore, if $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L-1},\overline{{\mathbf{W}}}_{L})=({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top},\cdots,{\mathbf{a}}_{L-1}{\mathbf{b}}_{L-1}^{\top},\overline{{\mathbf{w}}}^{\top}),$ where $\|{\mathbf{a}}_{l}\|_{2}=\|{\mathbf{b}}_{l}\|_{2}$, for all $l\in[L-1]$, then $|{\mathbf{a}}_{l}|=p^{1/4}|{\mathbf{b}}_{l+1}|$, for all $l\in[L-2]$, $|{\mathbf{a}}_{L-1}|=(p\hat{p})^{1/4}|\overline{{\mathbf{w}}}|$, $\|{\mathbf{a}}_{l}\|_{2}^{4}=p^{L-l}/\hat{p},$ for all $l\in[L-1],$ and $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/\hat{p}$, where $\hat{p}=p^{L-1}+p^{L-2}+\cdots+1.$
The above lemma implies $|{\mathbf{a}}_{l}|$ is parallel to $|{\mathbf{b}}_{l+1}|$. This is the first step towards establishing the alignment conditions stated in the above theorems. Now, if ${\mathbf{a}}_{l}$ and ${\mathbf{b}}_{l+1}$ are non-negative, which is true in certain cases, we get ${\mathbf{a}}_{l}$ is parallel to ${\mathbf{b}}_{l+1}.$ We similarly use additional conditions to get alignment results for other cases.
We next describe how to show ${\mathbf{b}}_{1}$ is a KKT point of the smaller optimization problem. For this, we focus on deep ReLU networks and assume that the alignment conditions are true, the proof of other activation functions follow a similar approach. From the KKT conditions, we know $\overline{{\mathbf{W}}}_{1}$ must satisfy